### v1

# Spectral Properties of Generalized Correlation Matrices

Abstract.
We introduce a family of coefficients that generalize the notion of correlation and explore their properties in the large dimensional multivariate case, showing that in the null case of uncorrelated variables, the spectrum of generalized correlation matrices is distributed according to an affine transformation of the Marčenko-Pastur law.
Key words and phrases:
random matrices, generalized correlation matrices, Marčenko-Pastur law
2020 Mathematics Subject Classification:
60B20, 62H20
1. Introduction
Pearson’s $\rho$, and Kendall’s $\tau$ are two of the most commonly used metrics for quantifying the strength of the relationship between random variables. Pearson’s $\rho$ assesses the linear relationships, making it particularly suitable for Gaussian data. In contrast, Kendall’s $\tau$ is a rank-based measure that offers greater robustness against heavy-tailed distributions. Understanding the unique characteristics of each coefficient is important for selecting the appropriate metric based on the application [7]. In this paper, however, we adopt a broader perspective by introducing a family of coefficients (see (1)) that generalizes the notion of correlation (it encompasses both [^1] Although closely related, Spearman’s rank correlation is not included in our generalization (see [9]). Pearson’s $\rho$ and Kendall’s $\tau$) and we explore their properties in the multivariate case, more specifically in the large dimensional regime for the null case of independent variables. This motivates potential applications in data analysis techniques such as feature extraction (see, e.g., [8] for applications to portfolio construction in finance).
Random Matrix Theory (RMT) offers powerful tools for analyzing the properties of correlation matrices, especially in high-dimensional settings. The empirical spectral distribution of a Hermitian matrix is the uniform distribution on its eigenvalues. One of the purposes of RMT is to understand the asymptotic behaviour, as the dimension grows, of such distributions. In our case, we are interested in the asymptotic behaviour of these distributions when the entries of the matrix are empirical correlation coefficients (of the new, general type introduced here at (1)) of uncorrelated variables.
The first significant result in this area appeared with the analysis of the empirical spectral distribution of large sample covariance matrices in the null case. More specifically, the seminal work of [10] proved that the empirical spectral distribution of the sample covariance matrix of a white noise converges to a Marčenko-Pastur (MP) law when both the data dimension and the sample size grow (at the same rate). This is striking given the true covariance matrix of the white noise is the identity matrix, with spectrum concentrated at 1. For Kendall’s $\tau$, [3] proved the convergence to an affine transformation of an MP law, elegantly using a variant of the Hoeffding decomposition.
Building on the work of [3], we show that the spectrum of generalized correlation matrices, with entries being generalized correlation coefficients, are also distributed according to an affine transformation of a MP law. This novel result, illustrated here by Figure 1, highlights the intricate relationship between MP distributions and sample correlation matrices of uncorrelated variables.
Figure 1. Histogram of the eigenvalues of $\mathcal{COR}^{\phi}$ when $d\!=\!2,\phi\!=\!\tanh,p\!=\!500$ and q=0.5q\!=\!0.5, with i.i.d. standard Gaussian data. The superimposed red line is the density function of the Marčenko-Pastur law $\mathbb{MP}_{q}^{t,1-t}$ with $t\approx 0.9$ (defined in Section 4.1), as predicted by our Theorem 2.1
2. Main result
For $d\geq 2$ and $n$ positive integers, $\mathcal{X}$ an arbitrary set, $\phi:\mathcal{X}^{d}\to\mathbb{R}$ antisymmetric [^2] A function $\phi:\mathcal{X}^{d}\to\mathbb{R}$ is said to be antisymmetric if for any $(x_{1},\ldots,x_{d})\in\mathcal{X}^{d}$ and $\pi$ permutation of $[d]$, we have $\phi(x_{\pi(1)},\ldots,x_{\pi(d)})=\operatorname{sgn}(\pi)\phi(x_{1},\ldots,x_{d})$ and $\mathsf{u},\mathsf{v}\in\mathcal{X}^{n}$, we define the $\phi$-covariance $\mathcal{COV}^{\phi}(\mathsf{u},\mathsf{v})$ of $\mathsf{u}$ and $\mathsf{v}$ as
$\mathcal{COV}^{\phi}(\mathsf{u},\mathsf{v})=\frac{1}{\binom{n}{d}}\sum_{(i_{1},\ldots,i_{d})\in[n]^{d}_{<}}\phi(\mathsf{u}_{i_{1}},\ldots,\mathsf{u}_{i_{d}})\phi(\mathsf{v}_{i_{1}},\ldots,\mathsf{v}_{i_{d}}),$ (1)
where $n]=\{1,\ldots,n\}$ and
$n]^{d}_{<}=\{(i_{1},\dots,i_{d})\in[n]^{d}\,;\,i_{1}<\cdots<i_{d}\}.$
When the denominator below is nonzero, the $\phi$-correlation $\mathcal{COR}^{\phi}(\mathsf{u},\mathsf{v})$ of $\mathsf{u}$ and $\mathsf{v}$ as
$\mathcal{COR}^{\phi}(\mathsf{u},\mathsf{v})=\frac{\mathcal{COV}^{\phi}(\mathsf{u},\mathsf{v})}{\sqrt{\mathcal{COV}^{\phi}(\mathsf{u},\mathsf{u})\mathcal{COV}^{\phi}(\mathsf{v},\mathsf{v})}}.$ (2)
Just as the covariance of two vectors extends to the notion of covariance matrix of a family of vectors, this notion extends to families of vectors: for $\mathsf{u}(1),\ldots,\mathsf{u}(p)\in\mathcal{X}^{n}$, we define the $p\times p$ matrix
$\mathcal{COV}^{\phi}(\mathsf{u}(1),\ldots,\mathsf{u}(p))=\begin{bmatrix}\mathcal{COV}^{\phi}(\mathsf{u}(k),\mathsf{u}(l))\end{bmatrix}_{1\leq k,l\leq p}$ (3)
and, when all involved denominators are nonzero,
$\mathcal{COR}^{\phi}(\mathsf{u}(1),\ldots,\mathsf{u}(p))=\begin{bmatrix}\mathcal{COR}^{\phi}(\mathsf{u}(k),\mathsf{u}(l))\end{bmatrix}_{1\leq k,l\leq p}$ (4)
The following theorem, illustrated by Figure 1, is the main result of this text.
Theorem 2.1.
Let $d\geq 2$ be an integer. For each $n$, let $p$ be a positive integer,
$\left(\mathsf{X}_{i}(k)\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq k\leq p\end{subarray}}$
be an array of i.i.d. random variables taking values in a measured space $\mathcal{S}$, and $\phi:\mathcal{S}^{d}\to\mathbb{R}$ be an antisymmetric measurable function (all, except $d$, implicitly depending on $n$). We introduce
$\mathcal{X}_{1},\ldots,\mathcal{X}_{d}$ (5)
some independent copies of $\mathsf{X}_{1}(1)$ and the function
$\overline{\phi}(x)=\operatorname{\mathbb{E}}\left[\phi(x,\mathcal{X}_{2},\ldots,\mathcal{X}_{d})\right]$ (6)
(so that the random variables $\mathcal{X}_{1},\ldots\mathcal{X}_{d}$ and the function $\overline{\phi}$ implicitly depend on $n$). Suppose that
  • the ratio $p/n$ tends to a finite limit $q\geq 0$ as $n\to\infty$,
  • $\operatorname{\mathbb{E}}\left[\phi(\mathcal{X}_{1},\ldots,\mathcal{X}_{d})^{2d+4}\right]$ is bounded by a finite constant independent of $n$,
  • the limits
    $\alpha=\lim_{n\to\infty}d\operatorname{\mathbb{E}}\left[\,\overline{\phi}(\mathcal{X}_{1})^{2}\,\right],\qquad\beta=\lim_{n\to\infty}\operatorname{\mathbb{E}}\left[\phi(\mathcal{X}_{1},\ldots,\mathcal{X}_{d})^{2}\right]$
    exist and $\beta>0$.
Then, as $n\to\infty$, for $\mathsf{X}(j)=(\mathsf{X}_{i}(j))_{1\leq i\leq n}$, the empirical spectral distribution [^3] The empirical spectral distribution of a Hermitian matrix is the uniform probability measure on its eigenvalues (counted with multiplicity). of
$\mathcal{COV}^{\phi}(\mathsf{X}(1),\ldots,\mathsf{X}(p))$
converges weakly in probability to the probability measure [^4] The probability distributions $\mathbb{MP}_{q}^{\alpha,\beta}$ are defined in Section 4.1. These are afffine transformations of Marčenko-Pastur laws. $\mathbb{MP}_{q}^{\alpha,\beta-\alpha}$ and the empirical spectral distribution of
$\mathcal{COR}^{\phi}(\mathsf{X}(1),\ldots,\mathsf{X}(p))$
converges weakly in probability to $\mathbb{MP}_{q}^{t,1-t}$ for $t=\frac{\alpha}{\beta}$.
Remark 2.2.
Note that as dimension $p$, the space $\mathcal{S}$, $\phi$, the distribution of $\mathsf{X}_{i}(k)$ might depend on $n$. The dependence in $n$ is dropped from notations for brevity. The only parameters that do not depend on $n$ are the fixed integer $d$, the limits $q$, $\alpha$ and $\beta$, and the implicit uniform bound on $\operatorname{\mathbb{E}}\left[\phi(\mathcal{X}_{1},\ldots,\mathcal{X}_{d})^{2d+4}\right]$.
Remark 2.3.
Note that the hypothesis that the limit of $\operatorname{\mathbb{E}}\left[\phi(\mathcal{X}_{1},\ldots,\mathcal{X}_{d})^{2}\right]$ is positive implies that with probability tending to one, none of the $\mathcal{COV}^{\phi}(\mathsf{X}(k),\mathsf{X}(k))$ (for $1\leq k\leq p$) is null (this can be seen as a consequence of Lemmas 3.1 and 3.4), so that
$\mathcal{COR}^{\phi}(\mathsf{X}(1),\ldots,\mathsf{X}(p))$
is actually well defined.
Remark 2.4.
When $d\geq 3$, by antisymmetry of $\phi$ and exchangeability of the $\mathcal{X}_{i}$’s, the function $\overline{\phi}$ defined at (6) is null. Hence, in this case, the statements of the theorem come down to say that the spectrum of $\mathcal{COV}^{\phi}(\mathsf{X}(1),\ldots,\mathsf{X}(p))$ (resp. of $\mathcal{COR}^{\phi}(\mathsf{X}(1),\ldots,\mathsf{X}(p))$) concentrates asymptotically in any neighborhood of $\beta$ (resp. of $1$), which is a non-trivial result. This opens the way to further investigations on the use of values of $d$ that are larger than $2$ to identify the null case of uncorrelated variables via the spectrum of such matrices, even at high ratios $p/n$.
Remark 2.5.
This generalization of correlation coefficients was inspired by [6], when $d=2$. In this case, a classical example of antisymmetric function is $\phi(x,y)=\psi(y-x)$ for $\psi:\mathbb{R}^{d}\to\mathbb{R}$ odd. Note that when $\psi(x)=x$, $\frac{1}{2}\mathcal{COV}^{\phi}(u,v)$ is exactly the standard unbiased estimator of covariance so that $\mathcal{COR}^{\phi}(u,v)$ is the Pearson’s $\rho$ correlation coefficient. Also, when $\psi(x)=\text{sgn}(x)$, $\mathcal{COR}^{\phi}(u,v)$ corresponds to the Kendall’s $\tau$ correlation coefficient. Thus, Theorem 2.1 recovers both the classical Marčenko-Pastur theorem and the result proved in [3] (with a relaxation of the hypotheses). For general $\mathcal{S}$, lots of examples for $d\geq 3$ can stem from the remark that for any antisymmetric function $\phi_{0}:\mathcal{S}^{2}\to\mathbb{R}$, the function $\phi:\mathcal{S}^{d}\to\mathbb{R}$ defined by
$\phi(x_{1},\ldots,x_{d})=\prod_{(i,j)\in[d]_{<}^{2}}\phi_{0}(x_{i},x_{j})$
is antisymmetric.
Remark 2.6.
The cornerstone of the proof of the theorem is Lemma 3.2, whose idea has been inspired by [3]. It can be summed-up as follows: in each entry
$\qquad\qquad\frac{1}{\binom{n}{d}}\sum_{(i_{1},\ldots,i_{d})\in[n]^{d}_{<}}\phi(\mathsf{X}_{i_{1}}(k),\ldots,\mathsf{X}_{i_{d}}(k))\phi(\mathsf{X}_{i_{1}}(l),\ldots,\mathsf{X}_{i_{d}}(l))\qquad\qquad\text{($1\leq k,l\leq p$)}$
of the matrix $\mathcal{COV}^{\phi}(\mathsf{X}(1),\ldots,\mathsf{X}(p))$, one can replace the average over $\{i_{2},\ldots,i_{d}\}$ by the corresponding theoretical expectation without changing the spectrum of the matrix at the macroscopic level. This is surprising given the whole Marčenko-Pastur theory relies on the fact that in an empirical covariance matrix, one cannot replace empirical means by their theoretical values without triggering a dramatic macroscopic change in the spectrum.
3. Proofs
3.1. Notation
Let $M^{\mathsf{T}}$ denote the transpose of a matrix $M$. For $u,v\in\mathbb{R}^{p}$, we set $u\otimes v=uv^{\mathsf{T}}$, where $u,v$ are considered as column vectors (so that $u\otimes v$ is a $p\times p$ matrix). Let us also introduce the operator $\mathbf{D}_{0}$ that maps any square matrix to the same matrix, where all diagonal entries have been set to zero and $\operatorname{Diag}$, the operator that maps any vector to the square diagonal matrix with diagonal entries the coordinates of the vector. We also use the notation $\operatorname{Diag}(M)=M-\mathbf{D}_{0}(M)$ for $M$ a square matrix. At last, we define the Frobenius norm $\left\|M\right\|_{\operatorname{F}}$ of a real matrix $M$ as follows
$\left\|M\right\|_{\operatorname{F}}=\sqrt{\operatorname{Tr}MM^{\mathsf{T}}}.$ (7)
3.2. Proof of Theorem 2.1
Let us prove Theorem 2.1 based on Lemmas 3.1, 3.2, 3.3, and 3.4, that will be proved below.
Remember that $\phi$ depends implicitly on $n$. For $\kappa>0$, we define $\phi_{\kappa}:\mathcal{S}^{d}\to\mathbb{R}$ as follows:
$\phi_{\kappa}(x_{1},\ldots,x_{d})=\operatorname{sign}(\phi(x_{1},\ldots,x_{d}))\min(|\phi(x_{1},\ldots,x_{d})|,n^{\kappa}).$
Then, $\overline{\phi}_{\kappa}$ is defined as $\overline{\phi}$: $\overline{\phi}_{\kappa}(x)=\operatorname{\mathbb{E}}\left[\phi_{\kappa}(x,\mathcal{X}_{2},\ldots,\mathcal{X}_{d})\right]$. Note that $\phi_{\kappa}$ is bounded by $n^{\kappa}$ and $\phi(x_{1},\ldots,x_{d})=\phi_{\kappa}(x_{1},\ldots,x_{d})$ when $|\phi(x_{1},\ldots,x_{d})|\leq n^{\kappa}$.
Lemma 3.1.
There is an exponent $\kappa\in(0,1/2)$ such that
$\mathbb{P}\left(\exists(i_{1},\ldots,i_{d},k)\in[n]^{d}\times[p]\text{ such that }|\phi(\mathsf{X}_{i_{1}}(k),\ldots,\mathsf{X}_{i_{d}}(k))|>n^{\kappa}\right)\underset{n\to\infty}{\longrightarrow}0$ (8)
and for any fixed integer $k\geq 3$,
$\frac{\operatorname{\mathbb{E}}\left[|\overline{\phi}_{\kappa}(\mathcal{X}_{1})|^{k}\right]}{n^{\frac{k}{2}-1}}\underset{n\to\infty}{\longrightarrow}0.$ (9)
By (8), the probability that
$\mathcal{COV}^{\phi}(\mathsf{X}(1),\ldots,\mathsf{X}(p))\neq\mathcal{COV}^{\phi_{\kappa}}(\mathsf{X}(1),\ldots,\mathsf{X}(p))$
or
$\mathcal{COR}^{\phi}(\mathsf{X}(1),\ldots,\mathsf{X}(p))\neq\mathcal{COR}^{\phi_{\kappa}}(\mathsf{X}(1),\ldots,\mathsf{X}(p))$
tends to zero as $n\to\infty$, so given we prove a convergence in probability, we can switch from $\phi$ to $\phi_{\kappa}$, i.e. , suppose that for a certain $\kappa\in(0,1/2)$,
$|\phi|\leq n^{\kappa}$ (10)
and for any $k\geq 3$,
$\frac{\operatorname{\mathbb{E}}\left[|\overline{\phi}(\mathcal{X}_{1})|^{k}\right]}{n^{\frac{k}{2}-1}}\underset{n\to\infty}{\longrightarrow}0.$ (11)
For $i\in[n]$, we define the vectors of $\mathbb{R}^{p}$
$\mathsf{X}_{i}(\bullet)$ $=$ $(\mathsf{X}_{i}(1),\ldots,\mathsf{X}_{i}(p))$
$\mathsf{U}_{i}$ $=$ $\overline{\phi}(\mathsf{X}_{i}(\bullet))$ (12)
where, in (12), $\overline{\phi}$ is applied entry-wise to each of the $p$ components.
The following equations are obvious from the fact that $\phi$ is antisymmetric and the $\left(\mathsf{X}_{i}(k)\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq k\leq p\end{subarray}}$ are i.i.d.  and will be used extensively in the following: for any $\mathbf{i}\in[n]^{d}_{<}$}, and any $k\in[p]$,
$\operatorname{\mathbb{E}}[\phi(\mathsf{X}_{i_{1}}(k),\ldots,\mathsf{X}_{i_{d}}(k))]=\operatorname{\mathbb{E}}[\overline{\phi}(\mathsf{X}_{i_{2}}(k))]=\cdots=\operatorname{\mathbb{E}}[\overline{\phi}(\mathsf{X}_{i_{d}}(k))]=0.$
Lemma 3.2.
Let us define the $p\times p$ matrix
$\Gamma=\frac{d}{n}\sum_{i=1}^{n}\mathsf{U}_{i}\otimes\mathsf{U}_{i}+\left(\operatorname{\mathbb{E}}\left[\phi(\mathcal{X}_{1},\ldots,\mathcal{X}_{d})^{2}\right]-d\operatorname{\mathbb{E}}\left[\,\overline{\phi}(\mathcal{X}_{1})^{2}\,\right]\right)\mathbf{I}_{p}.$ (13)
Then
$\operatorname{\mathbb{E}}\left[\left\|\mathcal{COV}^{\phi}(\mathsf{X}(1),\ldots,\mathsf{X}(p))-\Gamma\right\|_{\operatorname{F}}^{2}\right]=O(1).$
Lemma 3.3.
The empirical spectral law of the $p\times p$ matrix $\Gamma$ from (13) converges weakly in probability to $\mathbb{MP}_{q}^{\alpha,\beta-\alpha}$. Besides, we have $\operatorname{\mathbb{E}}[\left\|\Gamma\right\|_{\operatorname{F}}^{2}]=O(p).$
Now, the proof of the convergence of the empirical spectral law of
$\mathcal{COV}^{\phi}(\mathsf{X}(1),\ldots,\mathsf{X}(p))$
to $\mathbb{MP}_{q}^{\alpha,\beta-\alpha}$ is a direct application of Lemmas 3.2, 3.3 and 4.2.
Lemma 3.4.
[…]
$\mathbb{P}\left(\left|\mathcal{COV}^{\phi}(\mathsf{X}(k),\mathsf{X}(k))-\operatorname{\mathbb{E}}\left[\phi(\mathcal{X}_{1},\ldots,\mathcal{X}_{d})^{2}\right]\right|\geq r\right)\leq 2\mathrm{e}^{-2r^{2}/\sigma^{2}}$
for $\sigma^{2}=d^{2d+2}n^{2\kappa-1}$.
Let us now prove the convergence of the empirical spectral law of
$\mathcal{COR}^{\phi}(\mathsf{X}(1),\ldots,\mathsf{X}(p))$
to $\mathbb{MP}_{q}^{t,1-t}$ for $t=\frac{\alpha}{\beta}$. Note that
$\mathcal{COR}^{\phi}(\mathsf{X}(1),\ldots,\mathsf{X}(p))=D^{-1/2}\mathcal{COV}^{\phi}(\mathsf{X}(1),\ldots,\mathsf{X}(p))D^{-1/2}$
for $D$ the diagonal matrix defined by
$D=\begin{pmatrix}\mathcal{COV}^{\phi}(\mathsf{X}(1),\mathsf{X}(1))&&\\ &\ddots&\\ &&\mathcal{COV}^{\phi}(\mathsf{X}(p),\mathsf{X}(p))\end{pmatrix}$
By Lemma 3.4, there is a constant $C$ depending only on $d$ such that for any $k\in[p]$,
$\mathbb{P}\left(\left|\mathcal{COV}^{\phi}(\mathsf{X}(k),\mathsf{X}(k))-\operatorname{\mathbb{E}}\left[\phi(\mathcal{X}_{1},\ldots,\mathcal{X}_{d})^{2}\right]\right|\geq Cn^{\kappa/2-1/4}\right)\leq 2\mathrm{e}^{-n^{1/2-\kappa}},$
hence by the union bound,
$\mathbb{P}\left({{\left\|D-\operatorname{\mathbb{E}}\left[\phi(\mathcal{X}_{1},\ldots,\mathcal{X}_{d})^{2}\right]\mathbf{I}_{p}\right\|}}\geq Cn^{\kappa/2-1/4}\right)\leq 2p\mathrm{e}^{-n^{1/2-\kappa}},$
so that the hypotheses of Lemma 4.3 are satisfied for $M_{n}=\mathcal{COV}^{\phi}(\mathsf{X}(1),\ldots,\mathsf{X}(p))$, $\lambda=\beta$ and $f(x)=\mathbbm{1}_{x>0}x^{-1/2}$, which concludes the proof.
3.3. Proof of Lemma 3.1
By hypothesis, $\operatorname{\mathbb{E}}\left[\phi(\mathcal{X}_{1},\ldots,\mathcal{X}_{d})^{2d+4}\right]$ is bounded by a finite constant independent of $n$, so there is a constant $C$ such that uniformly in $\kappa>0$, $n$ and $(i_{1},\ldots,i_{d},k)\in[n]^{d}\times[p]$,
$\mathbb{P}\left(|\phi(\mathsf{X}_{i_{1}}(k),\ldots,\mathsf{X}_{i_{d}}(k))|>n^{\kappa}\right)\leq Cn^{-(2d+4)\kappa}$
Hence by the union bound, (8) is satisfied as soon as
$d+1-(2d+4)\kappa<0.$ (14)
Note that by the hypothesis on $\operatorname{\mathbb{E}}\left[\phi(\mathcal{X}_{1},\ldots,\mathcal{X}_{d})^{2d+4}\right]$ again, (9) is always satisfied for $k\in\{3,\ldots,2d+4\}$. For $k\geq 2d+5$, we have
$\operatorname{\mathbb{E}}\left[|\overline{\phi}_{\kappa}(\mathcal{X}_{1})|^{k}\right]\leq n^{k\kappa},$
so that (9) is satisfied as soon as $k\kappa<\frac{k}{2}-1$, i.e.
$\kappa<\frac{1}{2}-\frac{1}{k},$ (15)
which is satisfied for any $k\geq 2d+5$ as soon as it is satisfied for $k=2d+5$. Joining conditions (14) and (15), we get that the lemma is proved if we can find $\kappa\in(0,1/2)$ such that
$\frac{d+1}{2d+4}<\kappa<\frac{1}{2}-\frac{1}{2d+5}.$
Given
$\frac{d+1}{2d+4}+\frac{1}{2d+5}<\frac{d+1}{2d+4}+\frac{1}{2d+4}=\frac{1}{2},$
this concludes the proof.
3.4. Proof of Lemma 3.2
For $\mathbf{i}=(i_{1},\ldots,i_{d})\in[n]^{d}_{<}$, we define the vector of $\mathbb{R}^{p}$
$A_{\mathbf{i}}$ $=$ $\phi(\mathsf{X}_{i_{1}}(\bullet),\ldots,\mathsf{X}_{i_{d}}(\bullet)),$
where $\phi$ is applied entry-wise to each of the $p$ components.
Note that, by the very definition (1) of $\mathcal{COV}^{\phi}$, we have
$\mathcal{COV}^{\phi}(\mathsf{X}(1),\ldots,\mathsf{X}(p))=\frac{1}{\binom{n}{d}}\sum_{\mathbf{i}\in[n]^{d}_{<}}A_{\mathbf{i}}\otimes A_{\mathbf{i}}$
Note also the following relations between $A_{\mathbf{i}}$’s conditional expectations and the $\mathsf{U}_{i}$’s defined at (12):
$\operatorname{\mathbb{E}}\!\left[A_{\mathbf{i}}\mid\mathsf{X}_{i_{1}}(\bullet)\right]$ $=\mathsf{U}_{i_{1}},$ (16)
$\operatorname{\mathbb{E}}\!\left[A_{\mathbf{i}}\mid\mathsf{X}_{i_{l}}(\bullet)\right]$ $=-\,\mathsf{U}_{i_{l}},\qquad\forall\,l\in\{2,\ldots,d\}.$
Then, for $\mathbf{i}=(i_{1},\ldots,i_{d})\in[n]^{d}_{<}$, we define the vector of $\mathbb{R}^{p}$
$\overline{A}_{\mathbf{i}}$ $=$ $A_{\mathbf{i}}-\left(\mathsf{U}_{i_{1}}\!-\!\mathsf{U}_{i_{2}}\!-\!\dotsb\!-\!\mathsf{U}_{i_{d}}\right).$ (17)
Henceforth, we denote $\overline{U}_{\mathbf{i}}=\mathsf{U}_{i_{1}}\!-\!\mathsf{U}_{i_{2}}\!-\!\dotsb\!-\!\mathsf{U}_{i_{d}}$.
Then, we introduce the matrices $M^{(1)}_{\mathbf{i}},M^{(2)}_{\mathbf{i}},M^{(3)}_{\mathbf{i}}$ defined as follows
$M^{(1)}_{\mathbf{i}}$ $=\operatorname{Diag}\!\left(\phi^{2}(\mathsf{X}_{i_{1}}(\bullet),\ldots,\mathsf{X}_{i_{d}}(\bullet))\right)+\mathbf{D}_{0}\left(\overline{U}_{\mathbf{i}}\otimes\overline{U}_{\mathbf{i}}\right)$ (18)
$M^{(2)}_{\mathbf{i}}$ $=\mathbf{D}_{0}\left(\overline{A}_{\mathbf{i}}\otimes\overline{U}_{\mathbf{i}}\right)$
$M^{(3)}_{\mathbf{i}}$ $=\mathbf{D}_{0}\left(\overline{A}_{\mathbf{i}}\otimes\overline{A}_{\mathbf{i}}\right)$
where, in (18), $\phi^{2}$ is applied entry-wise to each of the $p$ components.
Then, using the identity, stemming from (17), that $A_{\mathbf{i}}=\overline{U}_{\mathbf{i}}+\overline{A}_{\mathbf{i}}$, we have the decomposition
$A_{\mathbf{i}}\otimes A_{\mathbf{i}}$ $=$ $\operatorname{Diag}(\phi^{2}(\mathsf{X}_{i_{1}}(\bullet),\ldots,\mathsf{X}_{i_{d}}(\bullet)))+\mathbf{D}_{0}(A_{\mathbf{i}}\otimes A_{\mathbf{i}})$
$=$ $\operatorname{Diag}(\phi^{2}(\mathsf{X}_{i_{1}}(\bullet),\ldots,\mathsf{X}_{i_{d}}(\bullet)))+\mathbf{D}_{0}\left(\overline{U}_{\mathbf{i}}\otimes\overline{U}_{\mathbf{i}}\right)+$
$\mathbf{D}_{0}\left(\overline{A}_{\mathbf{i}}\otimes\overline{U}_{\mathbf{i}}\right)+\mathbf{D}_{0}\left(\overline{U}_{\mathbf{i}}\otimes\overline{A}_{\mathbf{i}}\right)+$
$\mathbf{D}_{0}\left(\overline{A}_{\mathbf{i}}\otimes\overline{A}_{\mathbf{i}}\right)$
$=$ $M^{(1)}_{\mathbf{i}}+M^{(2)}_{\mathbf{i}}+(M^{(2)}_{\mathbf{i}})^{\mathsf{T}}+M^{(3)}_{\mathbf{i}}$
Then, the proof of Lemma 3.2 follows directly from Fact 3.5 and Fact 3.6 below, using the triangular inequality for the norm $\left\|\,\cdot\,\right\|_{\operatorname{F}}$ and Inequality (24).
Fact 3.5.
For $\Gamma$ as in (13), we have
$\operatorname{\mathbb{E}}\left[\left\|\frac{1}{\binom{n}{d}}\sum_{\mathbf{i}\in[n]^{d}_{<}}M^{(1)}_{\mathbf{i}}-\Gamma\right\|_{\operatorname{F}}^{2}\right]=O(1).$
Fact 3.6.
For any $l\in\{2,3\}$,
$\operatorname{\mathbb{E}}\left[\left\|\sum_{\mathbf{i}\in[n]^{d}_{<}}M^{(l)}_{\mathbf{i}}\right\|_{\operatorname{F}}^{2}\right]=O(n^{2(d-1)}p^{2})$
Proof of Fact 3.5. Using Lemma 4.1, we have
$\sum_{\mathbf{i}\in[n]^{d}_{<}}\overline{U}_{\mathbf{i}}\otimes\overline{U}_{\mathbf{i}}$ $=$ $\binom{n\!-\!1}{d\!-\!1}\sum_{i=1}^{n}\mathsf{U}_{i}\otimes\mathsf{U}_{i}+\sum_{\begin{subarray}{c}1\leq i,j\leq n\\ i\neq j\end{subarray}}\left(\binom{n\!-\!2}{d\!-\!2}-2\binom{n\!-\!\min(i,j)\!-\!1}{d\!-\!2}\right)\mathsf{U}_{i}\otimes\mathsf{U}_{j}$
Hence,
$\frac{1}{\binom{n}{d}}\sum_{\mathbf{i}\in[n]^{d}_{<}}M^{(1)}_{\mathbf{i}}-\Gamma$ $=\frac{1}{\binom{n}{d}}\sum_{\mathbf{i}\in[n]^{d}_{<}}\Big(\operatorname{Diag}(\phi^{2}(\mathsf{X}_{i_{1}}(\bullet),\ldots,\mathsf{X}_{i_{d}}(\bullet)))-\operatorname{\mathbb{E}}\left[\phi(\mathcal{X}_{1},\ldots,\mathcal{X}_{d})^{2}\right]\mathbf{I}_{p}\Big)$
$\quad-\frac{d}{n}\sum_{i=1}^{n}\Big(\operatorname{Diag}(\mathsf{U}_{i}\otimes\mathsf{U}_{i})-\operatorname{\mathbb{E}}\left[\,\overline{\phi}(\mathcal{X}_{1})^{2}\,\right]\mathbf{I}_{p}\Big)$
$\quad+\frac{1}{\binom{n}{d}}\mathbf{D}_{0}\,\sum_{\begin{subarray}{c}1\leq i,j\leq n\\ i\neq j\end{subarray}}\left(\binom{n\!-\!2}{d\!-\!2}-2\binom{n\!-\!\min(i,j)\!-\!1}{d\!-\!2}\right)\mathsf{U}_{i}\otimes\mathsf{U}_{j}$ (19)
The first term of the RHT of (19) is treated as follows, using (25):
$\operatorname{\mathbb{E}}\!\Big[\Big\|\frac{1}{\binom{n}{d}}\sum_{\mathbf{i}\in[n]^{d}_{<}}\!\Big(\operatorname{Diag}(\phi^{2}(\mathsf{X}_{i_{1}}(\bullet),\ldots,\mathsf{X}_{i_{d}}(\bullet)))-\operatorname{\mathbb{E}}\left[\phi(\mathcal{X}_{1},\ldots,\mathcal{X}_{d})^{2}\right]\mathbf{I}_{p}\Big)\Big\|_{F}^{\!2}\Big]=$
$\sum_{k=1}^{p}\operatorname{Var}\!\left(\frac{1}{\binom{n}{d}}\sum_{\mathbf{i}\in[n]^{d}_{<}}\phi^{2}(\mathsf{X}_{i_{1}}(k),\ldots,\mathsf{X}_{i_{d}}(k))\right)\;\leq\;\frac{2pd^{d}}{n}\,\operatorname{Var}\!\big(\phi(\mathcal{X}_{1},\ldots,\mathcal{X}_{d})^{2}\big),$ (20)
using the fact that $\binom{n}{d}\geq(n/d)^{d}$.
The second term of the RHT of (19) is treated as follows:
$\operatorname{\mathbb{E}}\left[\left\|\frac{d}{n}\sum_{i=1}^{n}\left(\operatorname{Diag}(\mathsf{U}_{i}\otimes\mathsf{U}_{i})-\operatorname{\mathbb{E}}\left[\,\overline{\phi}(\mathcal{X}_{1})^{2}\,\right]\mathbf{I}_{p}\right)\right\|_{\operatorname{F}}^{2}\right]=$
$\sum_{k=1}^{p}\operatorname{Var}\left(\frac{d}{n}\sum_{i=1}^{n}\overline{\phi}(\mathsf{X}_{i}(k))^{2}\right)\;=\;\frac{d^{2}p}{n}\operatorname{Var}(\overline{\phi}(\mathcal{X}_{1})^{2})$ (21)
The third term of the RHT of (19) is treated as follows:
$\operatorname{\mathbb{E}}\left[\left\|\frac{1}{\binom{n}{d}}\mathbf{D}_{0}\sum_{\begin{subarray}{c}1\leq i,j\leq n\\ i\neq j\end{subarray}}\left(\binom{n\!-\!2}{d\!-\!2}-2\binom{n\!-\!\min(i,j)\!-\!1}{d\!-\!2}\right)\mathsf{U}_{i}\otimes\mathsf{U}_{j}\right\|_{\operatorname{F}}^{2}\right]$
$=\frac{1}{\binom{n}{d}^{2}}\sum_{\begin{subarray}{c}1\leq k,l\leq p\\ k\neq l\end{subarray}}\operatorname{\mathbb{E}}\Bigg(\Bigg(\sum_{\begin{subarray}{c}1\leq i,j\leq n\\ i\neq j\end{subarray}}\left(\binom{n\!-\!2}{d\!-\!2}-2\binom{n\!-\!\min(i,j)\!-\!1}{d\!-\!2}\right)\overline{\phi}(\mathsf{X}_{i}(k))\overline{\phi}(\mathsf{X}_{j}(l))\Bigg)^{2}\Bigg)$
$\leq\frac{\binom{n\!-\!2}{d\!-\!2}^{2}}{\binom{n}{d}^{2}}\sum_{\begin{subarray}{c}1\leq k,l\leq p\\ k\neq l\end{subarray}}\operatorname{\mathbb{E}}\Bigg(\sum_{\begin{subarray}{c}1\leq i,j\leq n\\ i\neq j\end{subarray}}\overline{\phi}(\mathsf{X}_{i}(k))^{2}\,\overline{\phi}(\mathsf{X}_{j}(l))^{2}\Bigg)$
$\leq\frac{Kp^{2}}{n^{2}}\,\operatorname{\mathbb{E}}\big[\overline{\phi}(\mathcal{X}_{1})^{2}\big]^{2},$ (22)
for a constant $K>0$ depending only on $d$. Then, we conclude the proof of Fact 3.5 starting from (19), using the triangular inequality for the norm $\left\|\,\cdot\,\right\|_{\operatorname{F}}$ and (24), and treating the three terms on the RHT thanks to respectively (20), (21), and (22). $\square$
Proof of Fact 3.6. Let $l\in\{2,3\}$. First, for any $\mathbf{i}\in[n]^{d}_{<}$} and any $\mathbf{j}\in[n]^{d}_{<}$},
$\bigl|\{i_{1},\ldots,i_{d}\}\cap\{j_{1},\ldots,j_{d}\}\bigr|\leq 1\implies\operatorname{\mathbb{E}}\operatorname{Tr}(M^{(l)}_{\mathbf{i}}(M^{(l)}_{\mathbf{j}})^{\mathsf{T}})=0$
(this stems from Tower property, (16) and the fact that the $\mathsf{U}_{i}$’s are centered, as $\phi$ is antisymmetric). It follows that for
$\mathcal{C}_{d}=\Bigl\{(\mathbf{i},\mathbf{j})\in[n]^{d}_{<}\times[n]^{d}_{<}:\bigl|\{i_{1},\dots,i_{d}\}\cap\{j_{1},\dots,j_{d}\}\bigr|\geq 2\Bigr\},$
we have
$\operatorname{\mathbb{E}}\left[\left\|\sum_{\mathbf{i}\in[n]^{d}_{<}}M^{(l)}_{\mathbf{i}}\right\|_{\operatorname{F}}^{2}\right]$ $=$ $\sum_{(\mathbf{i},\mathbf{j})\in\mathcal{C}_{d}}\sum_{1\leq k,l\leq p}\operatorname{\mathbb{E}}\left[(M^{(l)}_{\mathbf{i}})_{k,l}(M^{(l)}_{\mathbf{j}})_{k,l}\right].$
This yields the conclusion, given that $\bigl|\mathcal{C}_{d}\bigr|=O(n^{2(d-1)})$ and that the second moments of the entries of $M^{(l)}_{\mathbf{i}}$ are dominated by the fourth moments of $\phi(\mathcal{X}_{1},\ldots,\mathcal{X}_{d})$, which are supposed to be bounded. $\square$
3.5. Proof of Lemma 3.3
When $q>0$, the convergence is a direct application of [4, Th. 3.2]. Indeed,
$\frac{1}{n}\sum_{i=1}^{n}\mathsf{U}_{i}\otimes\mathsf{U}_{i}=\frac{1}{n}MM^{\mathsf{T}}$
for
$M=[\overline{\phi}(\mathsf{X}_{i}(k))]_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq k\leq p\end{subarray}}$
and we have the bound (11) on the moments of the entries of $M$. The fact that $\operatorname{\mathbb{E}}[\left\|\Gamma\right\|_{\operatorname{F}}^{2}]=O(p)$ is for example a consequence of Proposition 4.4 (used with (24)).
For the case $q=0$, we simply use Proposition 4.4 with Lemma 4.2.
3.6. Proof of Lemma 3.4
The lemma is a direct application of Lemma 4.7 for
$f(x_{1},\ldots,x_{n})=\frac{1}{\binom{n}{d}}\sum_{\mathbf{i}\in[n]^{d}_{<}}\phi(x_{i_{1}},\ldots,x_{i_{d}})^{2}$
with the observation that given $|\phi|\leq n^{\kappa}$, the numbers $\sigma_{i}$ of (26) satisfy
$\sigma_{i}\leq\frac{dn^{d-1}n^{\kappa}}{\binom{n}{d}}\leq d^{d+1}n^{\kappa-1}$
(we used the fact that $\binom{n}{d}\geq(n/d)^{d}$), so that
$\sum_{i=1}^{n}\sigma_{i}^{2}\leq d^{2d+2}n^{2\kappa-1}.$
4. Appendix
4.1. Marčenko-Pastur distributions $\mathbb{MP}_{q}$ and $\mathbb{MP}_{q}^{\alpha,\beta}$
For $q\geq 0$, we define the Marčenko-Patsur distribution $\mathbb{MP}_{q}$ with parameter $q$ as follows:
$\mathbb{MP}_{q}=\begin{cases}\delta_{1}&\text{ if }q=0,\\ \frac{\sqrt{(\lambda_{+}-x)(x-\lambda_{-})}}{2\pi qx}\mathrm{d}x+\mathbbm{1}_{q>1}(1-q^{-1})\delta_{0}&\text{ if }q>0,\end{cases}$ (23)
for $\lambda_{\pm}=(1\pm\sqrt{q})^{2}$.
For $\alpha,\beta\in\mathbb{R}$, we set $\mathbb{MP}_{q}^{\alpha,\beta}$ to be the push-forward of $\mathbb{MP}_{q}$ by the map $x\mapsto\alpha x+\beta$.
4.2. A polynomial identity
The following identity is one of the keys of our proof.
Lemma 4.1.
Let $2\leq d\leq n$. In the space $\mathbb{C}[X_{1},Y_{1},\ldots,X_{n},Y_{n}]$ of polynomials in the indeterminates $X_{1},Y_{1},\ldots,X_{n},Y_{n}$, the following identity holds:
$\sum_{\mathbf{i}\in[n]^{d}_{<}}(X_{i_{1}}\!-\!X_{i_{2}}\!-\!\cdots\!-\!X_{i_{d}})(Y_{i_{1}}\!-\!Y_{i_{2}}\!-\!\cdots-\!Y_{i_{d}})$
$\qquad\;=\;\binom{n\!-\!1}{d\!-\!1}\sum_{i=1}^{n}X_{i}Y_{i}+\sum_{\begin{subarray}{c}1\leq i,j\leq n\\ i\neq j\end{subarray}}\left(\binom{n\!-\!2}{d\!-\!2}\!-\!2\binom{n\!-\!\min(i,j)\!-\!1}{d\!-\!2}\right)X_{i}Y_{j}.$
Proof.
[…]
$c_{r}=\begin{cases}+1,&r=1,\\ -1,&r=2,\ldots,m.\end{cases}$
Then, if $i$ is the smallest of the $d$-subset, $c_{\sf{rk}(i)}c_{\sf{rk}(j)}\!=\!-\!1$ and $c_{\sf{rk}(i)}c_{\sf{rk}(j)}\!=\!1$ otherwise. Therefore, for $i<j$, the term in front of $X_{i}Y_{j}$ is,
$(+1)\cdot\left(\binom{n\!-\!2}{d\!-\!2}\!-\!\binom{n\!-\!i\!-\!1}{d\!-\!2}\right)+(-1)\cdot\binom{n\!-\!i\!-\!1}{d\!-\!2}=\binom{n\!-\!2}{d\!-\!2}\!-\!2\binom{n\!-\!i\!-\!1}{d\!-\!2}$
The result follows by symmetry for general $i\neq j$. ∎
4.3. Approximation and Frobenius norm
The following lemma follows directly from [2, Cor. A.41].
Lemma 4.2.
For $n\geq 1$, let $M_{n},M_{n}^{\prime}$ be symmetric real matrices, both having size $p_{n}\times p_{n}$, with $p_{n}\to\infty$ as $n\to\infty$. Suppose that the empirical spectral distribution of $M_{n}$ converges weakly in probability to a non random limit probability measure $\mu$ and that, for $\left\|\,\cdot\,\right\|_{\operatorname{F}}$ the Frobenius norm defined at (7),
$\frac{\left\|M_{n}^{\prime}-M_{n}\right\|_{\operatorname{F}}^{2}}{p_{n}}.$
converges in probability to zero. Then the empirical spectral distribution of $M_{n}^{\prime}$ converges weakly in probability to $\mu$.
Let us denote by $\left\|\,\cdot\,\right\|$}} the operator norm associated with the canonical Euclidian norm. In the lemma below, $f(D_{n})$ has to be understood as the result of the functional calculus (i.e. $f$ is applied to the eigenvalues of $D_{n}$ and the eigenvectors remain unchanged).
Lemma 4.3.
For $n\geq 1$, let $M_{n},D_{n}$ be symmetric real matrices, both having size $p_{n}\times p_{n}$, with $p_{n}\to\infty$ as $n\to\infty$. Suppose that
  • (i)
    the empirical spectral distribution of $M_{n}$ converges weakly in probability to a non random limit probability measure $\mu$,
  • (ii)
    $\operatorname{\mathbb{E}}\left[\left\|M_{n}\right\|_{\operatorname{F}}^{2}\right]=O(p_{n})$,
  • (iii)
    there is $\lambda\in\mathbb{R}$ such that $\left\|D_{n}-\lambda\mathbf{I}_{p_{n}}\right\|$}} converges in probability to zero as $n\to\infty$.
Then for any measurable $f:\mathbb{R}\to\mathbb{R}$ continuous at $\lambda$, the empirical spectral distribution of $f(D_{n})M_{n}f(D_{n})$ converges weakly in probability to the push-forward of $\mu$ by the function $x\mapsto f(\lambda)^{2}x$.
Proof.
Note that (iii) simply means that any neighborhood of $\lambda$ will, with probability tending to one as $n\to\infty$, contain the whole spectrum of $D_{n}$. Hence by continuity of $f$ at $\lambda$, we can suppose that $f$ is the identity function. Then, by Lemma 4.2, one has to prove that
$\frac{\left\|D_{n}M_{n}D_{n}-\lambda^{2}M_{n}\right\|_{\operatorname{F}}^{2}}{p_{n}}.$
converges in probability to zero. For $\varepsilon_{n}=D_{n}-\lambda\mathbf{I}_{p_{n}}$, by (24), we have
$\left\|D_{n}M_{n}D_{n}-\lambda^{2}M_{n}\right\|_{\operatorname{F}}^{2}\leq 8|\lambda|\left\|\varepsilon_{n}M_{n}\right\|_{\operatorname{F}}^{2}+4\left\|\varepsilon_{n}^{2}M_{n}\right\|_{\operatorname{F}}^{2}.$
Then, one concludes using the non-commutative Hölder inequality [1, Appendix A.3] and the fact that convergence in probability to zero is robust to multiplication by $L^{1}$-bounded sequences. ∎
Now comes a statement that we need for the case $q=0$ and that we did not find in literature. Note that this elementary computation shows, without the need of the whole Marčenko-Pastur Theorem, that the empirical covariance matrix of a high-dimensional white noise cannot be close to the identity matrix when the sample size $n$ and the number of variables $p$ have the same order (more strongly, that the average distance from its eigenvalues to $1$ has order $\sqrt{p/n}$).
Proposition 4.4.
Let $M=\left[x_{k,t}\right]_{\begin{subarray}{c}1\leq k\leq p\\ 1\leq t\leq n\end{subarray}}$ be a $p\times n$ matrix of independent, centered, $L^{4}$ real random variables such that for all $k,t$, $\operatorname{\mathbb{E}}[x_{k,t}^{2}]=1.$ Then we have
$\operatorname{\mathbb{E}}\left[\left\|\frac{1}{n}MM^{\mathsf{T}}-\mathbf{I}_{p}\right\|_{\operatorname{F}}^{2}\right]\leq\frac{p}{n}\left(p+\max_{k,t}\operatorname{\mathbb{E}}[x_{k,t}^{4}]-2\right),$
with equality if all $x_{k,t}$’s have the same fourth moment.
Proof.
Given $\operatorname{\mathbb{E}}(MM^{\mathsf{T}})=\mathbf{I}_{p}$, the LHT reduces to $\frac{1}{n^{2}}\operatorname{\mathbb{E}}[\operatorname{Tr}(MM^{\mathsf{T}}MM^{\mathsf{T}})]-p.$ We have
$\operatorname{\mathbb{E}}[\operatorname{Tr}(MM^{\mathsf{T}}MM^{\mathsf{T}})]$ $=$ $\sum_{1\leq k,l\leq p}\sum_{1\leq t,s\leq n}\operatorname{\mathbb{E}}[x_{k,t}x_{l,t}x_{l,s}x_{k,s}]$
By independence, the terms in this sum corresponding to ($k\neq l$ and $s\neq t$) are null. The terms corresponding to ($k=l$ and $s=t$) have a total contribution upper-bounded by $np\max_{k,t}\operatorname{\mathbb{E}}[x_{k,t}^{4}]$ (with equality if the fourth moments coincide). The terms corresponding to ($k\neq l$ and $s=t$) (resp. to ($k=l$ and $s\neq t$)) have a total contribution equal to np(p1)np(p-1) (resp. pn(n1)pn(n-1)). One then concludes easily. ∎
The following identity, which can be proved by induction, is useful when dealing with the squared Frobenius norm of a sum.
Lemma 4.5.
For any $a_{1},\ldots,a_{k}\in\mathbb{R}$,
$(a_{1}+\cdots+a_{k})^{2}\leq 2^{k-1}(a_{1}^{2}+\cdots+a_{k}^{2}).$ (24)
4.4. Variance in a sparse graphical model
In the following lemma, whose proof is straightforward, the exponent 2d12d-1 (instead of the natural $2d$) is what has most of our convergences working.
Lemma 4.6.
Let
$\left(\mathsf{y}_{\mathbf{i}}\right)_{\mathbf{i}\in[n]^{d}_{<}}$
be an array of $L^{2}$ real random variables such that for all $\mathbf{i}=(i_{1},\ldots,i_{d})\in[n]^{d}_{<}$, $\mathbf{j}=(j_{1},\ldots,j_{d})\in[n]^{d}_{<}$,
$\{i_{1},\ldots,i_{d}\}\cap\{j_{1},\ldots,j_{d}\}=\emptyset\implies\operatorname{Cov}(\mathsf{y}_{\mathbf{i}},\mathsf{y}_{\mathbf{j}})=0.$
Then
$\operatorname{Var}\left(\sum_{\mathbf{i}\in[n]^{d}_{<}}\mathsf{y}_{\mathbf{i}}\right)\leq 2n^{2d-1}\max_{\mathbf{i}\in[n]^{d}_{<}}\operatorname{Var}(\mathsf{y}_{\mathbf{i}}).$ (25)
4.5. McDiarmid-Azuma-Hoeffding concentration inequality for bounded differences
The following lemma can be found, for example, in [11] or [5, Th. 6.2].
Lemma 4.7.
Let $X_{1},\ldots,X_{n}$ be random variances taking values in respectively $\mathcal{S}_{1},\ldots,\mathcal{S}_{n}$ and $f:\mathcal{S}_{1}\times\cdots\times\mathcal{S}_{n}\to\mathbb{R}$ measurable be such that for any $i\in[n]$,
$\sigma_{i}=\sup_{\begin{subarray}{c}x,x^{\prime}\in\mathcal{S}_{1}\times\cdots\times\mathcal{S}_{n}\\ \forall j\neq i,x_{j}=x^{\prime}_{j}\end{subarray}}|f(x^{\prime})-f(x)|$ (26)
is finite. Then, for any $r\geq 0$,
$\mathbb{P}\left(|f(X_{1},\ldots,X_{n})-\operatorname{\mathbb{E}}f(X_{1},\ldots,X_{n})|\geq r\right)\leq 2\mathrm{e}^{-2r^{2}/\sigma^{2}}$
for $\sigma^{2}$ defined by
$\sigma^{2}=\sigma_{1}^{2}+\cdots+\sigma_{n}^{2}.$
Acknowledgments
We thank Victor Le Coz, Matteo Smerlak and Jean-Philippe Bouchaud, who contributed to our research through fruitful discussions.
This research was primarily conducted within the Econophysics & Complex Systems Research Chair, under the aegis of the Fondation du Risque, the Fondation de l’Ecole Polytechnique, the Ecole Polytechnique and Capital Fund Management.